# Exact Solution of the Frustrated Potts Model with Next-Nearest-Neighbor Interactions in One Dimension: An AI-Aided Discovery

### v2

@@ Abstract
The one-dimensional $J_{1}$-$J_{2}$ $q$-state Potts model is solved exactly for arbitrary $q$ by introducing the maximally symmetric subspace (MSS) method to analytically block diagonalize the $q^{2}\times q^{2}$ transfer matrix to a simple $2\times 2$ matrix, based on using OpenAI’s latest reasoning model o3-mini-high to exactly solve the $q=3$ case. It is found that the model can be mapped to the 1D $q$-state Potts model with $J_{2}$ acting as the nearest-neighbor interaction and $J_{1}$ as an effective magnetic field, extending the previous proof for $q=2$, i.e., the Ising model. The exact results provide insights to outstanding physical problems such as the stacking of atomic or electronic orders in layered materials and the formation of a $T_{c}$-dome-shaped phase often seen in unconventional superconductors. This work is anticipated to fuel both the research in one-dimensional frustrated magnets for recently discovered finite-temperature application potentials and the fast moving topic area of AI for sciences.
@@
Finding novel phases and phase transitions is a central challenge in various research fields, including condensed matter physics, materials science, quantum information, and microelectronics [1]. Unusual phases abound in frustrated magnets [2], which are described typically by the Ising model [3] or the quantum Heisenberg model [4] with competing spin-spin interactions either in the form of an equilateral triangle or via competition between the nearest-neighbor (NN) interaction $J_{1}$ and next-nearest-neighbor (NNN) interaction $J_{2}$ [1].
[…]
@@
While the $J_{1}$-$J_{2}$ Ising model and Heisenberg model in one dimension [12, 13, 14, 15] and two dimension [16, 17, 18] have been extensively studied, only the 1D $J_{1}$-$J_{2}$ Ising model has been solved exactly by using the transfer matrix method [19]. Exact analytic solutions of the 1D $J_{1}$-$J_{2}$ Potts model also remain unknown; since the model with $q=3$ already exhibits a distinct ground-state phase behavior from that with $q=2$, i.e., the Ising model (Fig. 2[11], it is of fundamental importance to exactly solve the model for arbitrary $q$. The challenge arises from rapid increase in the order of the transfer matrix, which equals $q^{2}$. No wonder a $9\times 9$ matrix for $q=3$ is already hard to solve analytically and diagonalization of a $\left(10^{10^{10}}\right)^{2}\times\left(10^{10^{10}}\right)^{2}$ matrix for $q=10^{10^{10}}$ is simply beyond reach even numerically. Previous studies remarkably reduced the task to numerical calculations for an effective $q\times q$ matrix in the integer-$q$ formalism of the transfer matrix—and for an effective $2\times 2$ matrix in the continuous-$q$ formalism of the transfer matrix where physics is less transparent—however, short of analytic exact results [20]. Hence, an intuitive understanding of the rich phase behaviors in the 1D $J_{1}$-$J_{2}$ Potts model is still lacking.
@@
Two recent developments shed light on this long-standing problem. The first one is the analytic reduction of the $4\times 4$ transfer matrix for a decorated Ising ladder to an effective $2\times 2$ matrix using symmetry-based block diagonalization, leading to the discovery of spontaneous finite-temperature ultranarrow phase crossover (UNPC), which exponentially approaches the forbidden finite-temperature phase transition in 1D Ising models [21], and the subsequent discovery of in-field UNPC driven by exotic ice-fire states [22, 23, 24]. These findings point out the promising potentials of 1D frustrated magnets in finite-temperature applications; finding exact solutions for 1D frustrated Potts model could define a milestone in this important new direction. The second development is the derivation of an elegant equation—that determines the critical temperature of UNPC in site-decorated Ising models in an external magnetic field—by OpenAI’s latest reasoning model o3-mini-high at the first-ever 1000-Scientist AI Jam Session [24]. Hence, the author was inspired to prompt this AI reasoning model progressively to handle the transfer matrix in the integer-$q$ formalism for the $q=3$ case—despite quite a few errors in AI’s responses—and eventually have found a symmetry-based block diagonalization that can analytically reduce the $9\times 9$ transfer matrix of the 1D $J_{1}$-$J_{2}$ three-state Potts model to an effective $2\times 2$ matrix.
@@
For general $q$, the key symmetry is the full permutation symmetry of the $q$ Potts states. In other words, the Hamiltonian (and therefore the transfer matrix in the integer-$q$ formalism) is invariant under any permutation of the labels $\{1,2,3,...,q\}$; its symmetry group is $\mathcal{S}_{q}$. Although the AI failed to go further but warned that the number of permutations increases dramatically as $q$ increases, the exact results for the $q=2$ and $3$ cases—especially the point that both arrive at an effective $2\times 2$ matrix—stimulated the author to realize that since only the largest eigenvalue ($\lambda$) of the transfer matrix matters in the thermodynamic limit, the task is reduced to identify the symmetry-separated subspace that contains $\lambda$. Then, the author found that this subspace is spanned by two maximally symmetric vectors because all the transfer matrix elements are positive, resulting in an analytic $2\times 2$ matrix. Therefore, the surprisingly simple exact solution of the 1D $J_{1}$-$J_{2}$ Potts model has been obtained for arbitrary $q$.
@@
To construct the transfer matrix, we use both the overlapping-pairs formulation for Equation (1) with one spin per unit cell to get $\mathbb{T}$ and the equivalent zigzag ladder version of the model with two spins per unit cell [Fig. 1(b) and Equation (S1) in End Matter A] to obtain $\mathbb{T}^{\prime}$; they satisfy $\mathbb{T}^{\prime}=\mathbb{T}^{2}$. In the thermodynamic limit $N\to\infty$, the partition function $Z=\mathrm{Tr}\,e^{-\beta H}=\lambda^{2N}$, where $\lambda$ is the largest eigenvalue of the transfer matrix $\mathbb{T}$. The free energy per spin is given by
@@
The resulting transformation matrix $U_{q2}=\{|\psi_{1}\rangle,|\psi_{2}\rangle\}$ is a $q^{2}\times 2$ matrix that projects the $q^{2}\times q^{2}$ transfer matrix $\mathbb{T}$ to the $2\times 2$ block $\mathbb{T}_{2}=U_{q2}^{T}\;\mathbb{T}\;U_{q2}$, which is decoupled from the rest thanks to different symmetry and given by
@@
Equation (4) is the same as the projected transfer matrix for the 1D $q$-state Potts model with $J_{2}$ acting as the NN interaction and $J_{1}$ as an effective magnetic field [25]
$H_{\mathrm{eff}}=-J_{2}\sum_{i=1}^{2N}\delta({\sigma_{i},\sigma_{i+1}})-J_{1}\sum_{i=1}^{2N}\delta({\sigma_{i},1}),$ (6)
in the maximally symmetric subspace of its $q\times q$ transfer matrix, expanded by the following two $q\times 1$ vectors:
@@
This mapping was proven for the 1D $J_{1}$-$J_{2}$ Ising model ($q=2$) using the substitution of $\tau_{i}=\sigma_{i}\sigma_{i+1}$ and $\tau_{i}\tau_{i+1}=\sigma_{i}\sigma_{i+2}$ where $\sigma_{i}=\pm 1$ is the Ising spin values [12, 13]. It is now proven to apply to arbitrary $q$.
Figure 2: Phase diagrams for $q=2,3,4,\mathrm{and}~{}10^{6}$. Left panels: Density plots of the normalized entropy $2S(J_{1},T)/\ln q$ in the $J_{1}-T$ plane. Right panels: The temperature dependence of $2S(J_{1},T)/\ln q$ for selected $J_{1}$ values around the critical points. $-J_{2}=1$ is set as the energy unit.
The simplicity of Equation (4) provides an intuitive understanding of rich phase behaviors in the 1D $J_{1}$-$J_{2}$ Potts model. We are interested in frustrated cases where $J_{2}<0$ is antiferromagnetic and have set $-J_{2}=1$ as the energy unit from now on. The $q$-dependent phase diagrams given by the normalized entropy $2S(J_{1},T)/\ln q$ are shown in Fig. 2. A $T_{c}$-dome-like phase emerges at low temperature for small $q$ and disappears for large $q$. The phase diagrams given by $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle$ and $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle$ are shown in Fig. S1 in End Matter.
To get insights to these rich phase diagrams, we start with analyzing ground-state phase behaviors. At $T=0$, the 1D $J_{1}$-$J_{2}$ Potts model for all $q$ values have three phases separated by two critical points (CPs), determined by the relative magnitude of $u$, $w$, $v$ in Equation (4). The right-hand-side phase for $J_{1}>2$ (where $u$ dominates) is always “ferromagnetic” with $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=1$ and $E=-J_{1}-J_{2}$; more precisely, it is a mixture of $q$ ferromagnetic states, i.e., $\frac{1}{\sqrt{q}}(|111...\rangle+|222...\rangle+\cdots+|qqq...\rangle$. The “ferromagnetic transition” CP at $J_{1}=2$ has a residual entropy per spin
@@
$q=2$: The $J_{1}>0$ and $J_{1}<0$ regions are symmetry related by flipping the spins on one sublattice. Consequently, the left-hand-side phase for $J_{1}<-2$ is “antiferromagnetic” with $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=0$, $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=1$, $E=-J_{2}$. The two CPs are located at $J_{1}=\pm 2$. The CP at $J_{1}=-2$ has the same residual entropy as that at $J_{1}=2$ given by Equation (8), i.e., $\ln(\frac{1+\sqrt{5}}{2})$, which is indeed the same as the magnetic-field-induced CP in the 1D Ising model with only NN interactions [26]. The middle phase for $-2<J_{1}<2$ is a dimerized state, i.e., a pair of NN spins have the same variable but differ from the last pair, forming the patterns of …11221122… and its three degenerate shifts, leading to $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=1/2$, $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=0$, and $E=-J_{1}/2$. None of the three phases has a macroscopic degeneracy.
[…]
Figure 3: The $q\geq 3$ dependence of the zero-temperature normalized entropy $2S(J_{1},0)/\ln q$ for four different regions of $J_{1}$: $S=\ln(q-2)$ for $J_{1}<0$, $S=\ln(q-1)$ for $J_{1}=0$, $S=\ln\sqrt{q-1}$ for $20<J_{1}<$ 2, and $S=\ln[(1+\sqrt{4q-3})/2]$ for $J_{1}=2$. The $q=2$ case, which need be specially treated for different $J_{1}$ regions (see text), was added for comparison. $J_{2}=-1$.
@@
The $T_{c}$ dome is a key phenomenon in unconventional superconductivity in cuprates, iron-based superconductors, twisted bilayer graphene, etc. [27]. It has been explained in the picture of (i) a preformed order with phase coherence gradually building up [28] or (ii) two competing phases [29]. The present $q$-dependent emerging and vanishing of a dome-like structure, controlled by the relative strength of the residual entropies of the phase’s two CPs, provides another possibility for forming a dome-shaped phase.
In summary, the 1D $J_{1}$-$J_{2}$ $q$-state Potts model has been solved exactly by discovering that the largest eigenvalue of the $q^{2}\times q^{2}$ transfer matrix lives in a $2\times 2$ maximally symmetric subspace, and found to be equivalent to the 1D $q$-state Potts model with $J_{2}$ acting as the NN interaction and $J_{1}$ as the magnetic field. The model’s ground state are found to feature three phases separated by two critical points for all $q$ values. The relative strength of the two critical points’ residual entropies is large and small for small and large $q$, respectively, leading to the emerging and vanishing of a dome-shaped randomly dimerized phase for small and large $q$, respectively, providing a new mechanism for forming a dome-shaped phase. Moreover, the exact results provide insight to the stacking problems of atomic or electronic orders in layered materials. The maximally symmetric subspace method is anticipated to offer a significant advantage over previous numerical methods as the range of interactions increases, which will be addressed in subsequent publications. These discoveries were based on using OpenAI’s latest reasoning model o3-mini-high to exactly solve the $q=3$ case, which embodies a new paradigm of data- and value-driven discovery, inspiring scientists to evaluate and connect insights drawn from the vast—though imperfect—information provided by AI.

### v3

# Exact solution of the frustrated Potts model with next-nearest-neighbor interactions in one dimension via AI bootstrapping

@@ Abstract
The one-dimensional (1D) $J_{1}$-$J_{2}$ $q$-state Potts model is solved exactly for arbitrary $q$ by analytically block-diagonalizing the original $q^{2}\times q^{2}$ transfer matrix into a simple $2\times 2$ maximally symmetric subspace, based on using OpenAI’s reasoning model o3-mini-high to exactly solve the $q=3$ case. Furthermore, by matching relevant subspaces, we map the Potts model onto a simpler effective 1D $q$-state Potts model, where $J_{2}$ acts as the nearest-neighbor interaction and $J_{1}$ as an effective magnetic field, nontrivially generalizing a 56-year-old theorem previously limited to the simplest case ($q=2$, the Ising model). Our exact results provide insights to phenomena such as atomic or electronic order stacking in layered materials and the emergence of dome-shaped phases in complex phase diagrams. This work is anticipated to fuel both research in 1D frustrated magnets for recently discovered finite-temperature application potentials and the fast moving topic area of AI in science.
DOI: 10.1103/y5vc-3t6q
I Introduction
@@
Finding novel phases and phase transitions is a central challenge in various research fields, including condensed matter physics, materials science, quantum information, and microelectronics [1]. Unusual phases abound in frustrated magnets [2, 3], which are described typically by the Ising model or the quantum Heisenberg model with competing nearest-neighbor (NN) interaction $J_{1}$ and next-nearest-neighbor (NNN) interaction $J_{2}$ [1, 2, 3, 4].
Another fundamental model of statistical mechanics is the $q$-state Potts model [5, 6, 7, 8], which is a generalization of the Ising model ($q=2$) and can serve as a useful intermediary to study the transition from discrete (Ising) to continuous (Heisenberg) symmetry. In particular, the one-dimensional (1D) $J_{1}$-$J_{2}$ Potts model could be relevant to problems ranging from the out-of-plane stacking of atomic or electronic orders in layered materials, such as charge stripe ordering in La1.67Sr0.33NiO4 [9], the Star-of-David charge-density wave in 1TT-TaS2 [10], and spin spiral ordering in the Weyl semimetal EuAuSb [11]—where in-plane ordering is governed primarily by detailed quantum-mechanical effects, owing to in-plane interactions that are far stronger than those between planes—to a time series with multiple choices at every time step such as table tennis training drill designs.
@@
While the $J_{1}$-$J_{2}$ Ising and Heisenberg models in one dimension [12, 13, 14, 15] and two dimensions [16, 17, 18] have been extensively studied, only the 1D $J_{1}$-$J_{2}$ Ising model has been solved exactly by using the transfer matrix method [19]. Exact solutions of the 1D $J_{1}$-$J_{2}$ Potts model remain unknown as well. The challenge arises from rapid increase in the order of the transfer matrix, which equals $q^{2}$. No wonder a $9\times 9$ matrix for $q=3$ is already hard to solve analytically and diagonalization of a $\left(10^{10^{10}}\right)^{2}\times\left(10^{10^{10}}\right)^{2}$ matrix for $q=10^{10^{10}}$ is simply beyond reach even numerically. Previous studies remarkably reduced the task to numerical calculations for an effective $q\times q$ matrix in the integer-$q$ formalism of the transfer matrix—and for an effective $2\times 2$ matrix in the continuous-$q$ formalism of the transfer matrix where physics is less transparent—however, short of analytic exact results [20]. Hence, an intuitive understanding of the rich phase behaviors in the 1D $J_{1}$-$J_{2}$ Potts model is still lacking. Since the model with $q=3$ already exhibits a distinct ground-state phase behavior from that with $q=2$ (cf. Fig. 2), it is of fundamental importance to exactly solve the model for arbitrary $q$.
@@
Two recent developments shed light on this long-standing problem. The first one is the analytic reduction of the $4\times 4$ transfer matrix for a decorated Ising ladder to an effective $2\times 2$ matrix using symmetry-based block diagonalization, leading to the discovery of spontaneous finite-temperature ultranarrow phase crossover (UNPC), which exponentially approaches the forbidden finite-temperature phase transition in 1D Ising models [21], and the subsequent discovery of in-field UNPC driven by exotic ice-fire states [22, 23, 24]. These findings point out the promising potentials of 1D frustrated magnets in finite-temperature applications; finding exact solutions for 1D frustrated Potts model could define a milestone in this important new direction. The second development is the derivation of an elegant equation, which determines the critical temperature of UNPC in decorated Ising models in an external magnetic field, by OpenAI’s reasoning model o3-mini-high at the first-ever 1000-Scientist AI Jam Session [24]. Hence, the author was inspired to prompt this AI reasoning model progressively and reflectively to handle the transfer matrix in the integer-$q$ formalism for the $q=3$ case—despite quite a few errors in AI’s responses—and eventually have found a symmetry-based block diagonalization that can analytically reduce the $9\times 9$ transfer matrix of the 1D $J_{1}$-$J_{2}$ three-state Potts model to an effective $2\times 2$ matrix (see Supplemental Material I [25]).
@@
For general $q$, the key symmetry is the full permutation symmetry of the $q$ Potts states. In other words, the Hamiltonian (and therefore the transfer matrix in the integer-$q$ formalism) is invariant under any permutation of the labels $\{1,2,3,...,q\}$; its symmetry group is $\mathcal{S}_{q}$. Although the AI failed to go further but warned that the number of permutations increases dramatically as $q$ increases, the analytic results for the $q=2$ and $3$ cases—especially those that both arrive at an effective $2\times 2$ matrix—stimulated the author to realize that since only the largest eigenvalue ($\lambda$) of the transfer matrix matters in the thermodynamic limit, the task is reduced to identify the symmetry-separated subspace that contains $\lambda$. Then, the author found that this subspace is spanned by two maximally symmetric vectors, according to the Perron–Frobenius theorem for positive matrices [26, 27] (see Supplemental Material II). The resulting $2\times 2$ matrix is easy to solve; therefore, the surprisingly simple exact solution of the 1D $J_{1}$-$J_{2}$ Potts model has been obtained for arbitrary $q$.
The maximally symmetric subspace (MSS) method is in spirit similar to the effective Hamiltonian method for studying strongly correlated electronic materials [28, 29, 30]—with the effective model parameters depending on $J_{1}$, $J_{2}$, $q$, and temperature (see Supplemental Material III.A [25]). Furthermore, a rare exact mapping between two canonical Hamiltonians naturally emerges within the MSS method, namely the 1D $J_{1}$-$J_{2}$ $q$-state Potts model can be mapped onto a simpler 1D $q$-state Potts model, where $J_{2}$ acts as the NN interaction and $J_{1}$ as an effective magnetic field. This not only nontrivially generalizes a 56-year-old theorem previously limited to $q=2$ [12] but also profoundly bridges two research areas focused on spontaneous behaviors and external field effects, respectively, thereby demonstrating relevant-subspace-matching as a broadly applicable theoretical approach to tackling complex problems.
II Model and Solution
@@
To construct the transfer matrix, we use both the overlapping-pairs formulation for Equation (1) with one spin per unit cell to get $\mathbb{T}$ and the equivalent zigzag-ladder version of the model with two spins per unit cell [Fig. 1(b)] to obtain $\mathbb{T}^{\prime}$; they satisfy $\mathbb{T}^{\prime}=\mathbb{T}^{2}$ (see Supplemental Material I [25]). In the thermodynamic limit $N\to\infty$, the partition function $Z=\mathrm{Tr}\,e^{-\beta H}=\lambda^{2N}$, where $\lambda$ is the largest eigenvalue of the transfer matrix $\mathbb{T}$. The free energy per spin is given by
@@
The resulting transformation matrix $U_{q2}=\{|\psi_{1}\rangle,|\psi_{2}\rangle\}$ is a $q^{2}\times 2$ matrix that projects the $q^{2}\times q^{2}$ transfer matrix $\mathbb{T}$ to the $2\times 2$ block $\mathbb{T}_{2}=U_{q2}^{\top}\,\mathbb{T}\,U_{q2}$, which is decoupled from the rest thanks to different symmetry and given by
@@
III Mapping
The 1D $J_{1}$-$J_{2}$ $q$-state Potts model can be mapped onto a simpler $q$-state Potts model, where $J_{2}$ acts as a NN interaction and $J_{1}$ appears as an effective magnetic field:
$H_{\mathrm{eff}}=-J_{2}\sum_{i=1}^{2N}\delta({\sigma_{i},\sigma_{i+1}})-J_{1}\sum_{i=1}^{2N}\delta({\sigma_{i},1}).$ (6)
Proof: By the Perron–Frobenius theorem [26], the largest eigenvalue of the $q\times q$ transfer matrix of Equation (S5) lives in the MSS spanned by the following two $q\times 1$ vectors:
@@ III Mapping
The resulting $2\times 2$ matrix [31] is exactly the same as Equation (S6). ∎
This mapping was previously demonstrated for the simplest case by exploiting a special property unique to $q=2$ [12, 13, 14]. Here, its nontrivial generalization to arbitrary $q$ was achieved by matching the MSS of the two $q$-state Potts models. It implies that an antiferromagnetic Potts model in an external magnetic field—which is more tunable in practice—is as fundamental as a Potts model spontaneously frustrated by competing interactions (see Supplemental Material III.B [25]).
With hindsight, if the mapping for arbitrary $q$ had been conjectured, it could have been verified by numerical calculations. Thus, this conjecture would have provided a hint to the general analytical solution.
IV Results and Discussion
The simplicity of Equation (S6) provides an intuitive understanding of rich phase behaviors in the 1D $J_{1}$-$J_{2}$ Potts model. We are interested in frustrated cases where $J_{2}<0$ is antiferromagnetic and have set $J_{2}=1$ as the energy unit from now on. The $q$-dependent phase diagrams given by the normalized entropy $2S(J_{1},T)/\ln q$ are shown in Fig. 2. A $T_{c}$-dome-like phase emerges at low temperature for small $q$ and disappears for large $q$. The phase diagrams given by $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle$ and $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle$ are shown in Fig. 4 in Appendix B.
To get insights to these rich phase diagrams, we start with analyzing ground-state phase behaviors. At $T=0$, the 1D $J_{1}$-$J_{2}$ Potts model for all $q$ values have three phases separated by two critical points (CPs), determined by the relative magnitude of $u$, $w$, $v$ in Equation (S6). The right-hand-side phase for $J_{1}>2$ (where $u$ dominates) is always “ferromagnetic” with $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=1$ and $E=-J_{1}-J_{2}$; more precisely, it is a mixture of $q$ ferromagnetic states, i.e., $\frac{1}{\sqrt{q}}(|111...\rangle+|222...\rangle+\cdots+|qqq$$\rangle.$ The “ferromagnetic transition” CP at $J_{1}=2$ has a residual entropy per spin
@@ IV Results and Discussion
Figure 2: Phase diagrams for $q=2,3,4,\mathrm{and}~10^{6}$. Left panels: Density plots of the normalized entropy $2S(J_{1},T)/\ln q$ in the $J_{1}-T$ plane. Right panels: The temperature dependence of $2S(J_{1},T)/\ln q$ for selected $J_{1}$ values around the critical points. $J_{2}=1$ is set as the energy unit.
$q=2$: The $J_{1}>0$ and $J_{1}<0$ regions are symmetry related by flipping the spins on one sublattice. Consequently, the left-hand-side phase for $J_{1}<-2$ is “antiferromagnetic” with $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=0$, $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=1$, $E=-J_{2}$. The two CPs are located at $J_{1}=\pm 2$. The CP at $J_{1}=-2$ has the same residual entropy as that at $J_{1}=2$ given by Equation (8), i.e., $\ln(\frac{1+\sqrt{5}}{2})$, which is indeed the same as the magnetic-field-induced CP in the 1D Ising model with only NN interactions [32]. The middle phase for $2<J_{1}<2$ is a dimerized state, i.e., a pair of NN spins have the same variable but differ from the last pair, forming the patterns of …11221122… and its three degenerate shifts, leading to $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=1/2$, $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=0$, and $E=-J_{1}/2$. None of the three phases has a macroscopic degeneracy.
$q\geq 3$: The middle phase for $0<J_{1}<2$ (where $w$ dominates) is a randomly dimerized state (RDS) with $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=1/2$, $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=0$, and $E=-J_{1}/2$, where like in the dimerized phase for $q=2$, a pair of NN spins have the same variable but differ from the last pair, so it has q1q-1 choices, yielding a residual entropy of $\ln(q-1)$ per two spins or $\ln\sqrt{q-1}$ per spin. This phase could be relevant to the stacking problem of the Star-of-David charge-density wave in 1TT-TaS2 [10]. The left-hand-side phase for $J_{1}<0$ (where $v$ dominates) is a paramagnet with $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=0$ and $E=0$, where any spin differs from the two last spins, so it has q2q-2 choices, yielding a residual entropy of $\ln(q-2)$ per spin; notably, $q=3$ is special since the residual entropy is zero, forming the patterns of …123123… and its five degenerate states by permutations of {1, 2, 3}, reminiscent of the ABC stacking pattern of the charge stripe order in La1.67Sr0.33NiO4 [9] and the 120 spin spiral order in the Weyl semimetal EuAuSb [11]. In the CP at $J_{1}=0$, any spin differs from its last next-nearest neighbor, so it has q1q-1 choices, yielding a residual entropy of $\ln(q-1)$ per spin; again, $q=3$ is special because its residual entropies at both CPs are equal to $\ln 2$.
Figure 3: The $q\geq 3$ dependence of the zero-temperature normalized entropy $2S(J_{1},0)/\ln q$ for four different regions of $J_{1}$: $S=\ln(q-2)$ for $J_{1}<0$, $S=\ln(q-1)$ for $J_{1}=0$, $S=\ln\sqrt{q-1}$ for $0<J_{1}<2$, and $S=\ln[(1+\sqrt{4q-3})/2]$ for $J_{1}=2$. The $q=2$ case, which need be specially treated for different $J_{1}$ regions (see text), was added for comparison. $J_{2}=-1$.
@@
The $q$-dependent emerging and vanishing of a dome-like structure, controlled by the relative strength of the residual entropies of the phase’s two CPs, provides one more mechanism for forming a dome-shaped phase, a key phenomenon in complex materials, in addition to the two well-known pictures of (i) a preformed order with phase coherence gradually building up [33] and (ii) competing phases [34, 35].
V Summary
The 1D $J_{1}$-$J_{2}$ $q$-state Potts model has been solved exactly by discovering that the largest eigenvalue of the $q^{2}\times q^{2}$ transfer matrix lives in a $2\times 2$ maximally symmetric subspace, which is equivalent to that of the simpler 1D $q$-state Potts model with $J_{2}$ acting as the NN interaction and $J_{1}$ as the magnetic field. The model’s ground state are found to feature three phases separated by two critical points for all $q$ values. The relative strength of the two critical points’ residual entropies is large and small for small and large $q$, respectively, leading to the emerging and vanishing of a dome-shaped randomly dimerized phase for small and large $q$, respectively, providing a new mechanism for forming a dome-shaped phase and insights to the stacking problems of atomic or electronic orders in layered materials. These discoveries were based on using OpenAI’s latest reasoning model o3-mini-high to exactly solve the $q=3$ case, which embodies a new paradigm of data- and value-driven discovery, inspiring scientists to evaluate and connect insights drawn from the vast—though imperfect—information provided by AI.
Note added. We recently employed the MSS method to demonstrate that the ”half ice, half fire” driven ultranarrow phase crossover persists in 1D decorated Potts models for all $q>2$, exhibiting a $T_{0}$-dome structure in the phase diagram that is absent in the $q=2$ case [36].
Figure 4: Phase diagrams in the $J_{1}-T$ plane for $q=2,3,4,\mathrm{and}~10^{6}$ given by the density plots of $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle$ (top panels) and $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle$ (bottom panels). $J_{2}=1$ is set as the energy unit.
Acknowledgment